At a gas station it takes Ms. Ullian 2 minutes to put 5 gallons of gas in her car. How long will it take her to fill her tank that holds 12 gallons?
Answer by rfer(16322)

2/5=x/12
5x=24
x=24/5
x=4.8 minutes